I need to make a program that when asked the number of minutes you sleep. ie 437. It will tell you 17:00 to 00:17, 18:01 to 01:18. Basically the times when the first and last digits are backwards and 437 minutes apart. I am having a hard time setting up the loops and if statements. How to I go about this? How many while or for loops will I need? Please help.
1
Expert's answer
2012-03-30T11:43:42-0400
# the function counts the number of minutes between two times
# Input:

# t1, t2 - strings, represeting time in the form
"HH:MM",
# e.g. "15:32"
# Output:
# d - distance
between these times in minutes.
# It is assumed that t1 is
before t2.
# Even if t1="22:00" and t2="00:45",
#
then the distance is 45 + 2*60 = 165

def mins (t1, t2):
# compte the
minutes from 00:00 till t1 and t2
m1 = int(t1[0:2])*60 + int(t1[3:5])
m2 =
int(t2[0:2])*60 + int(t2[3:5])

# if m2 > m1 then we should return
m2-m1,
# otherwise, we should return
# m2 + (number of minutes in 24
hours=60*24-1440) - m1

d = m2-m1;
if (d<0):
d=d+1440
return
d




# tests

# "22:00", "00:45" => 165
print
mins("22:00", "00:45")

# "12:45", "13:00" => 15
print
mins("12:45", "13:00")